# The Identity Problem in the special affine group of $\mathbb{Z}^{2}$

### v5

@@ 5. Virtual solvability
Let $H$ be a virtually solvable subgroup of $\mathsf{SL}(2,\mathbb{Z})$. Then $H$ is either finite, or it contains a finite index subgroup $H^{\prime}$ isomorphic to $\mathbb{Z}$.
@@
By Theorem 3, let $F\leq\mathsf{SL}(2,\mathbb{Z})$ be a free subgroup over two generators, such that the index $[\mathsf{SL}(2,\mathbb{Z}):F]$ is finite. Then $[H:F\cap H]\leq[\mathsf{SL}(2,\mathbb{Z}):F]$, so the group $H^{\prime}\coloneqq F\cap H$ is of finite index in $H$. But $H^{\prime}$ is a subgroup of the free group $F$, so it must be free by Theorem 3. On the other hand, $H^{\prime}$ is a subgroup of the virtually solvable group $H$, so $H^{\prime}$ is virtually solvable. Since $H^{\prime}\leq\mathsf{SL}(2,\mathbb{Z})$ is virtually solvable and free, it must be abelian by Theorem 5. Therefore either $H^{\prime}\cong\mathbb{Z}$ or $H^{\prime}=\{I\}$; in the second case, $H$ is finite.
@@
By Lemma 11 and 13, we can already prove the first part of Proposition 7. In particular, by Lemma 11, $H$ is either finite or contains a finite index subgroup isomorphic to $\mathbb{Z}$. If $H$ is finite, then either it is trivial or it contains a non-trivial torsion element. If $H$ contains a finite index subgroup isomorphic to $\mathbb{Z}$, then by Lemma 13, either it contains a torsion element, or it is isomorphic to $\mathbb{Z}$. Therefore, we have proved that exactly one of the following holds.
@@
    $H$ contains a torsion element.
@@
which is the lattice of integer points in the $\mathbb{Q}$-linear space spanned by $\Lambda$. A basis of $\overline{\Lambda}$ is effectively computable in PTIME using the Hermite Normal Form [Bre11, Chapter 14]. It is then decidable in PTIME whether $\overline{\Lambda}=\Lambda$, again using the Hermite Normal Form.
@@ 5.1. $H$ is trivial
Note that the matrices in $\mathcal{G}$ commute. Therefore by Lemma 1, $\langle\mathcal{G}\rangle$ is a group if and only if Equation (6) has a solution $(n_{1},\ldots,n_{K})\in\mathbb{Z}_{>0}^{K}$. By the homogeneity of Equation (6), it has a solution in $\mathbb{Z}_{>0}^{K}$ if and only if it has a solution in $\mathbb{Q}_{>0}^{K}$. This is decidable in PTIME by linear programming.
@@ 5.2. $H$ contains a torsion element
Here, $I-T$ is invertible because $T\in\mathsf{SL}(2,\mathbb{Z})$ is non-trivial torsion, so the eigenvalues of $T$ are all different from one. We conclude that $(I,\boldsymbol{0})$ can be represented by a full-image word over the alphabet $\mathcal{G}$. Hence, $\langle\mathcal{G}\rangle$ is a group by Lemma 1.
@@ 5.3. $H$ is generated by a twisted inversion $A$
By direct computation, $(A,\boldsymbol{a})^{2}\cdot(A^{-1},\boldsymbol{b})^{3}\cdot(A,\boldsymbol{a})^{2}\cdot(A^{-1},\boldsymbol{b})=(I,A^{-1}(A+I)^{2}\boldsymbol{a}+(A+I)^{2}\boldsymbol{b}).$ But since $A$ is a twisted inversion, we have $(A+I)^{2}=0$. Therefore, $(A,\boldsymbol{a})^{2}\cdot(A^{-1},\boldsymbol{b})^{3}\cdot(A,\boldsymbol{a})^{2}\cdot(A^{-1},\boldsymbol{b})=(I,\boldsymbol{0})$. We conclude that $(I,\boldsymbol{0})$ can be represented as a full-image word over $\mathcal{G}$. Hence, $\langle\mathcal{G}\rangle$ is a group by Lemma 1.
@@ 5.4. $H$ is generated by a shear $A$
Since $A$ is a shear, it is triangularizable over $\mathbb{Q}$. Since $A_{1},\ldots,A_{K}$ all commute, they all have the same eigenvalue one, and are simultaneously triangularizable. Let $P$ be a matrix (with entries in $\mathbb{Q}$) such that $P^{-1}A_{i}P$ are all of the form $\begin{pmatrix}1&\lambda_{i}\\ 0&1\\ \end{pmatrix}$ with $\lambda_{i}\in\mathbb{Q}$. Then, we have the following conjugation:
@@ 5.5. $H$ is generated by an inverting scale $A$
    $\frac{a_{+}}{1-\lambda}-\frac{a_{-}}{1-\lambda^{-1}}=0$, but $\frac{b_{+}}{1-\lambda^{-1}}-\frac{b_{-}}{1-\lambda}\neq 0$. When $m$ is even, $\boldsymbol{w}_{m}$ is a positive multiple of $\left(\frac{b_{+}}{1-\lambda^{-1}}-\frac{b_{-}}{1-\lambda}\right)\boldsymbol{x}_{W}$, and when $n$ is odd, $\boldsymbol{w}_{n}$ is a negative multiple of $\left(\frac{b_{+}}{1-\lambda^{-1}}-\frac{b_{-}}{1-\lambda}\right)\boldsymbol{x}_{W}$. Therefore, there exists positive real numbers $r_{1},r_{2}$ such that $r_{1}\boldsymbol{w}_{m}+r_{2}\boldsymbol{w}_{n}=\boldsymbol{0}$. Since $\boldsymbol{w}_{m},\boldsymbol{w}_{n}$ have integer entries, there even exist positive integers $n_{1},n_{2}$ such that $n_{1}\boldsymbol{w}_{m}+n_{2}\boldsymbol{w}_{n}=\boldsymbol{0}$. Therefore $(I,\boldsymbol{0})=(I,\boldsymbol{w}_{m})^{n_{1}}(I,\boldsymbol{w}_{n})^{n_{2}}$ can be represented by a full-image word.
@@
    $\frac{b_{+}}{1-\lambda^{-1}}-\frac{b_{-}}{1-\lambda}=0$, but $\frac{a_{+}}{1-\lambda}-\frac{a_{-}}{1-\lambda^{-1}}\neq 0$. This case is the exact symmetry of the previous case.
@@
    Both $\frac{a_{+}}{1-\lambda}-\frac{a_{-}}{1-\lambda^{-1}}$ and $\frac{b_{+}}{1-\lambda^{-1}}-\frac{b_{-}}{1-\lambda}$ are non-zero. In this case, consider the vectors $\boldsymbol{v}_{2m}$, $\boldsymbol{v}_{2m+1}$, $\boldsymbol{w}_{2m}$, $\boldsymbol{w}_{2m+1}$ when $m$ tends to infinity. Since $\lambda<-1$, we have $\lim_{m\rightarrow\infty}(1-\lambda^{2m})=-\infty$, $\lim_{m\rightarrow\infty}(1-\lambda^{2m+1})=+\infty$, $\lim_{m\rightarrow\infty}(1-\lambda^{-2m})=1$, $\lim_{m\rightarrow\infty}(1-\lambda^{-2m-1})=1$. Therefore, the direction of $\boldsymbol{v}_{2m}$ tends towards $-\left(\frac{a_{+}}{1-\lambda}-\frac{a_{-}}{1-\lambda^{-1}}\right)\boldsymbol{x}_{V}$, the direction of $\boldsymbol{v}_{2m+1}$ tends towards $\left(\frac{a_{+}}{1-\lambda}-\frac{a_{-}}{1-\lambda^{-1}}\right)\boldsymbol{x}_{V}$, the direction of $\boldsymbol{w}_{2m}$ tends towards $\left(\frac{b_{+}}{1-\lambda^{-1}}-\frac{b_{-}}{1-\lambda}\right)\boldsymbol{x}_{W}$, the direction of $\boldsymbol{w}_{2m+1}$ tends towards $-\left(\frac{b_{+}}{1-\lambda^{-1}}-\frac{b_{-}}{1-\lambda}\right)\boldsymbol{x}_{W}$. Hence, when $m$ is large enough, the four vectors $\boldsymbol{v}_{2m}$, $\boldsymbol{v}_{2m+1}$, $\boldsymbol{w}_{2m}$, $\boldsymbol{w}_{2m+1}$ in $\mathbb{Z}^{2}$ generate $\mathbb{Q}^{2}$ as a $\mathbb{Q}_{\geq 0}$-cone. Hence, there exist positive integers $t_{1},\ldots,t_{4}$ such that
    $(I,\boldsymbol{v}_{2m})^{t_{1}}\cdots(I,\boldsymbol{w}_{2m+1})^{t_{4}}=(I,\boldsymbol{0}).$
@@
In all cases, $(I,\boldsymbol{0})$ can be represented by a full-image word, so $\langle\mathcal{G}\rangle$ is a group.
@@ 5.6. $H$ is generated by a positive scale $A$
These are the forms of $A_{i}$ and $\boldsymbol{a}_{i}$ under the new basis $(\boldsymbol{x}_{V},\boldsymbol{x}_{W})$. In particular we have $A\boldsymbol{x}_{V}=\lambda\boldsymbol{x}_{V},A\boldsymbol{x}_{W}=\lambda^{-1}\boldsymbol{x}_{W}$.
@@
$d_{ija}\coloneqq-\frac{a_{i}}{1-\lambda^{n_{i}}}+\frac{a_{j}}{1-\lambda^{-n_{j}}},\;\;d_{ijb}\coloneqq\frac{b_{i}}{1-\lambda^{n_{i}}}-\frac{b_{j}}{1-\lambda^{-n_{j}}}.$
@@
$({A^{\prime}_{i}},\boldsymbol{a}^{\prime}_{i})^{n_{j}}({A^{\prime}_{j}},\boldsymbol{a}^{\prime}_{j})^{n_{i}}=\left(I,(1-\lambda^{-n_{i}n_{j}})\cdot\left(\lambda^{n_{i}n_{j}}d_{ija},d_{ijb}\right)^{\top}\right).$
@@
$({A^{\prime}_{k}},\boldsymbol{a}^{\prime}_{k})^{n_{k}}=\left(I,\boldsymbol{e}_{k}\right).$
Denote by $\mathcal{G}^{\prime}$ the new alphabet $\{(A^{\prime}_{1},\boldsymbol{a}^{\prime}_{1}),\ldots,(A^{\prime}_{K},\boldsymbol{a}^{\prime}_{K})\}$. The following lemma shows that the rays $\boldsymbol{v}\mathbb{R}_{>0}$ such that $(I,\boldsymbol{v})\in\langle\mathcal{G}^{\prime}\rangle$ can “fill up” the cells $\operatorname{Cell}(\boldsymbol{d}_{ij}),(i,j)\in J_{+}\times J_{-}$.
@@
Let $(i,j)$ be a pair in $J_{+}\times J_{-}$ and $(x,y)^{\top}\in\mathbb{R}^{2}$ be a vector in $\operatorname{Cell}(\boldsymbol{d}_{ij})$. If $\lambda>1$, then there exist elements $(I,\boldsymbol{v})$ and $(I,\boldsymbol{w})$ in $\langle\mathcal{G}^{\prime}\rangle$, where the vector $(x,y)^{\top}$ can be written as $r_{1}\boldsymbol{v}+r_{2}\boldsymbol{w}$ for some $r_{1},r_{2}\in\mathbb{R}_{>0}$. Furthermore, $(I,\boldsymbol{v})$ and $(I,\boldsymbol{w})$ are represented by words over the alphabet $\mathcal{G}^{\prime}$, such that the letters $(A^{\prime}_{i},\boldsymbol{a}^{\prime}_{i})$ and $(A^{\prime}_{j},\boldsymbol{a}^{\prime}_{j})$ both occur. See Figure 10 for an illustration.
@@
For any $p\in\mathbb{Z}_{>0}$, denote
@@
If $\operatorname{Cell}(\boldsymbol{d}_{ij})$ is of dimension one or zero, then either $d_{ija}=0$ or $d_{ijb}=0$. In both cases, $(x,y)^{\top},\boldsymbol{v}_{p},\boldsymbol{w}_{p}$ are linearly dependant and have the same direction, so $(x,y)^{\top}$ can be written as $r_{1}\boldsymbol{v}_{p}+r_{2}\boldsymbol{w}_{p}$ for some $r_{1},r_{2}\in\mathbb{R}_{>0}$.
If $\operatorname{Cell}(\boldsymbol{d}_{ij})$ is of dimension two, then $d_{ija}\neq 0$ and $d_{ijb}\neq 0$. Let $p$ be large enough so that $\frac{\lambda^{pn_{i}n_{j}}d_{ija}}{d_{ijb}}>\frac{x}{y}>\frac{d_{ija}}{\lambda^{pn_{i}n_{j}}d_{ijb}}$. Then $(x,y)^{\top}$ can be written as $r_{1}\boldsymbol{v}_{p}+r_{2}\boldsymbol{w}_{p}$ for some $r_{1},r_{2}\in\mathbb{R}_{>0}$.
@@
Suppose $J_{+}$ and $J_{-}$ are non-empty. Let $k$ be an element in $J_{0}$, and $(x,y)^{\top}\in\mathbb{R}^{2}$ be a vector in $\operatorname{Cell}(\boldsymbol{e}_{k})$. If $\lambda>1$, then there exist elements $(I,\boldsymbol{v})$ and $(I,\boldsymbol{w})$ in $\langle\mathcal{G}^{\prime}\rangle$, such that the vector $(x,y)^{\top}$ can be written as $r_{1}\boldsymbol{v}+r_{2}\boldsymbol{w}$ for some $r_{1},r_{2}\in\mathbb{R}_{>0}$. Furthermore, $(I,\boldsymbol{v})$ and $(I,\boldsymbol{w})$ are represented as words over the alphabet $\mathcal{G}^{\prime}$, such that the letter $(A^{\prime}_{k},\boldsymbol{a}^{\prime}_{k})$ occurs.
@@
If $\operatorname{Cell}(\boldsymbol{e}_{k})$ has dimension one or zero, then take $p=0$ and the statement is trivial. Suppose that $\operatorname{Cell}(\boldsymbol{e}_{k})$ has dimension two, then we have $d_{ija}\neq 0$ and $d_{ijb}\neq 0$. Fix a large enough $p$ so that $\frac{\lambda^{pn_{i}n_{j}}a_{k}}{\lambda^{-pn_{i}n_{j}}b_{k}}>\frac{x}{y}>\frac{\lambda^{-pn_{i}n_{j}}a_{k}}{\lambda^{pn_{i}n_{j}}b_{k}}$. Then $(x,y)^{\top}$ can be written as $r_{1}(\lambda^{pn_{i}n_{j}}a_{k},\lambda^{-pn_{i}n_{j}}b_{k})^{\top}+r_{2}(\lambda^{-pn_{i}n_{j}}a_{k},\lambda^{pn_{i}n_{j}}b_{k})^{\top}$ for some $r_{1},r_{2}\in\mathbb{R}_{>0}$. When $q$ tends towards infinity, the direction of the vectors $\boldsymbol{v}_{pq},\boldsymbol{w}_{pq}$ tends respectively to $(\lambda^{pn_{i}n_{j}}a_{k},\lambda^{-pn_{i}n_{j}}b_{k})^{\top}$ and $(\lambda^{-pn_{i}n_{j}}a_{k},\lambda^{pn_{i}n_{j}}b_{k})^{\top}$. Therefore, for large enough $q$, the vector $(x,y)^{\top}$ can be written as $r^{\prime}_{1}\boldsymbol{v}_{pq}+r^{\prime}_{2}\boldsymbol{w}_{pq}$ for some $r^{\prime}_{1},r^{\prime}_{2}\in\mathbb{R}_{>0}$.
@@
  1. 1.
    If $t_{1}t_{M}=1$. Suppose $t_{1}=t_{M}=1$, the case where $t_{1}=t_{M}=-1$ can be done analogously.
@@
    where $\boldsymbol{y}$ and $\boldsymbol{y}^{\prime}$ can be written as a $\mathbb{R}_{>0}$-linear combination of elements in $\mathcal{S}$. Therefore $\boldsymbol{x}=\boldsymbol{y}+\boldsymbol{y}^{\prime}$ also satisfies this claim.
  2. 2.
    If $t_{1}t_{M}=-1$. In this case, $C_{2}\cdots C_{M-1}=C_{1}C_{2}\cdots C_{M}$ $=I$. Define
@@
    Then the product of $w^{\prime}$ is of the form $(I,\boldsymbol{x}^{\prime})$, where $\boldsymbol{x}^{\prime}$ is either $\boldsymbol{0}$ (when $w^{\prime}$ is empty), or $\boldsymbol{x}^{\prime}$ is a $\mathbb{R}_{>0}$-combination of elements in $\mathcal{S}$ by the induction hypothesis. Hence

### v6

@@ 5. Virtual solvability
Let $H$ be a finitely generated virtually solvable subgroup of $\mathsf{SL}(2,\mathbb{Z})$. Then $H$ is either finite, or it contains a finite index subgroup $H^{\prime}$ isomorphic to $\mathbb{Z}$.
@@
By Theorem 3, let $F\leq\mathsf{SL}(2,\mathbb{Z})$ be a free subgroup over two generators, such that the index $[\mathsf{SL}(2,\mathbb{Z}):F]$ is finite. Then $[H:F\cap H]\leq[\mathsf{SL}(2,\mathbb{Z}):F]$, so the group $H^{\prime}\coloneqq F\cap H$ is of finite index in $H$ and hence finitely generated. But $H^{\prime}$ is a subgroup of the free group $F$, so it must be free by Theorem 3. On the other hand, $H^{\prime}$ is a subgroup of the virtually solvable group $H$, so $H^{\prime}$ is virtually solvable. Since $H^{\prime}\leq\mathsf{SL}(2,\mathbb{Z})$ is virtually solvable and free, it must be abelian by Theorem 5. Therefore either $H^{\prime}\cong\mathbb{Z}$ or $H^{\prime}=\{I\}$; in the second case, $H$ is finite.
@@
By Lemma 11 and 13, we can already prove the first part of Proposition 7. In particular, by Lemma 11, $H$ is either finite or contains a finite index subgroup isomorphic to $\mathbb{Z}$. If $H$ is finite, then either it is trivial or it contains a non-trivial torsion element. If $H$ contains a finite index subgroup isomorphic to $\mathbb{Z}$, then by Lemma 13, either it contains a non-trivial torsion element, or it is isomorphic to $\mathbb{Z}$. Therefore, we have proved that exactly one of the following holds.
@@
    $H$ contains a non-trivial torsion element.
@@
which is the lattice of integer points in the $\mathbb{Q}$-linear space spanned by $\Lambda$. A basis of $\overline{\Lambda}$ can be computed in PTIME the following way. One can compute in PTIME a $\mathbb{Q}$-basis of the orthogonal space of $\sum_{i=1}^{m}\mathbb{Q}\boldsymbol{\ell}_{i}$, and thus write $\overline{\Lambda}$ as the integer solution set of a system of homogeneous linear equations of polynomial size. One can multiply the coefficients of these homogeneous linear equations by their common denominator, and suppose all the coefficients to be integers. From this system, a basis of $\overline{\Lambda}$ is computable in PTIME using row reduction. It is then decidable in PTIME whether $\overline{\Lambda}=\Lambda$, by checking whether each element in the basis of $\overline{\Lambda}$ belongs to $\Lambda$.
@@ 5.1. $H$ is trivial
Note that the matrices in $\mathcal{G}$ commute. Therefore by Lemma 1, $\langle\mathcal{G}\rangle$ is a group if and only if Equation (6) has a solution $(n_{1},\ldots,n_{K})\in\mathbb{Z}_{>0}^{K}$. By the homogeneity of Equation (6), it has a solution in $\mathbb{Z}_{>0}^{K}$ if and only if it has a solution in $\mathbb{Q}_{>0}^{K}$. This is decidable in PTIME by linear programming [Kha79].
@@ 5.2. $H$ contains a torsion element
Here, $I-T$ is invertible because $T\in\mathsf{SL}(2,\mathbb{Z})$ is non-trivial torsion, so the eigenvalues of $T$ are all different from one ($T$ must be in case (i) or (ii) in the classification of elements in $\mathsf{SL}(2,\mathbb{Z})$, Section 2). We conclude that $(I,\boldsymbol{0})$ can be represented by a full-image word over the alphabet $\mathcal{G}$. Hence, $\langle\mathcal{G}\rangle$ is a group by Lemma 1.
@@ 5.3. $H$ is generated by a twisted inversion $A$
By direct computation, $(A,\boldsymbol{a})^{2}\cdot(A^{-1},\boldsymbol{b})^{3}\cdot(A,\boldsymbol{a})^{2}\cdot(A^{-1},\boldsymbol{b})=(I,A^{-1}(A+I)^{2}\boldsymbol{a}+(A+I)^{2}\boldsymbol{b}).$ But since $A$ is a twisted inversion, we have $(A+I)^{2}=0$, so the claim is proved. We conclude that $(I,\boldsymbol{0})$ can be represented as a full-image word over $\mathcal{G}$. Hence, $\langle\mathcal{G}\rangle$ is a group by Lemma 1.
@@ 5.4. $H$ is generated by a shear $A$
Since $A$ is a shear, it is triangularizable over $\mathbb{Q}$. Since $A_{1},\ldots,A_{K}$ are all powers of $A$, they all have eigenvalue one and share the same 1-dimensional invariant subspace, and are therefore simultaneously triangularizable. Let $P$ be a matrix (with entries in $\mathbb{Q}$) such that $P^{-1}A_{i}P$ are all of the form $\begin{pmatrix}1&\lambda_{i}\\ 0&1\\ \end{pmatrix}$ with $\lambda_{i}\in\mathbb{Q}$. Then, we have the following conjugation:
@@ 5.5. $H$ is generated by an inverting scale $A$
    $\frac{a_{+}}{1-\lambda}-\frac{a_{-}}{1-\lambda^{-1}}=0$, but $\frac{b_{+}}{1-\lambda^{-1}}-\frac{b_{-}}{1-\lambda}\neq 0$. When $m$ is even, $\boldsymbol{w}_{m}$ is a positive multiple of $\left(\frac{b_{+}}{1-\lambda^{-1}}-\frac{b_{-}}{1-\lambda}\right)\boldsymbol{x}_{W}$, and when $n$ is odd, $\boldsymbol{w}_{n}$ is a negative multiple of $\left(\frac{b_{+}}{1-\lambda^{-1}}-\frac{b_{-}}{1-\lambda}\right)\boldsymbol{x}_{W}$. Therefore, there exist positive real numbers $r_{1},r_{2}$ such that $r_{1}\boldsymbol{w}_{m}+r_{2}\boldsymbol{w}_{n}=\boldsymbol{0}$. Since $\boldsymbol{w}_{m},\boldsymbol{w}_{n}$ have integer entries, there even exist positive integers $n_{1},n_{2}$ such that $n_{1}\boldsymbol{w}_{m}+n_{2}\boldsymbol{w}_{n}=\boldsymbol{0}$. Therefore $(I,\boldsymbol{0})=(I,\boldsymbol{w}_{m})^{n_{1}}(I,\boldsymbol{w}_{n})^{n_{2}}$ can be represented by a full-image word.
@@
    $\frac{b_{+}}{1-\lambda^{-1}}-\frac{b_{-}}{1-\lambda}=0$, but $\frac{a_{+}}{1-\lambda}-\frac{a_{-}}{1-\lambda^{-1}}\neq 0$. By symmetry, this case follows from the previous one.
@@
    Both $\frac{a_{+}}{1-\lambda}-\frac{a_{-}}{1-\lambda^{-1}}$ and $\frac{b_{+}}{1-\lambda^{-1}}-\frac{b_{-}}{1-\lambda}$ are non-zero. In this case, consider the vectors $\boldsymbol{v}_{2m}$, $\boldsymbol{v}_{2m+1}$, $\boldsymbol{w}_{2m}$, $\boldsymbol{w}_{2m+1}$ when $m$ tends to infinity. Since $\lambda<-1$, we have $\lim_{m\rightarrow\infty}(1-\lambda^{2m})=-\infty$, $\lim_{m\rightarrow\infty}(1-\lambda^{2m+1})=+\infty$, $\lim_{m\rightarrow\infty}(1-\lambda^{-2m})=1$, $\lim_{m\rightarrow\infty}(1-\lambda^{-2m-1})=1$. Therefore, the direction of $\boldsymbol{v}_{2m}$ tends towards $-\left(\frac{a_{+}}{1-\lambda}-\frac{a_{-}}{1-\lambda^{-1}}\right)\boldsymbol{x}_{V}$, the direction of $\boldsymbol{v}_{2m+1}$ tends towards $\left(\frac{a_{+}}{1-\lambda}-\frac{a_{-}}{1-\lambda^{-1}}\right)\boldsymbol{x}_{V}$, the direction of $\boldsymbol{w}_{2m}$ tends towards $\left(\frac{b_{+}}{1-\lambda^{-1}}-\frac{b_{-}}{1-\lambda}\right)\boldsymbol{x}_{W}$, the direction of $\boldsymbol{w}_{2m+1}$ tends towards $-\left(\frac{b_{+}}{1-\lambda^{-1}}-\frac{b_{-}}{1-\lambda}\right)\boldsymbol{x}_{W}$. Hence, when $m$ is large enough, the four vectors $\boldsymbol{v}_{2m}$, $\boldsymbol{v}_{2m+1}$, $\boldsymbol{w}_{2m}$, $\boldsymbol{w}_{2m+1}$ in $\mathbb{Z}^{2}$ generate $\mathbb{Q}^{2}$ as a $\mathbb{Q}_{\geq 0}$-cone. Hence, there exist positive integers $t_{1},t_{2},t_{3},t_{4}$ such that
    $(I,\boldsymbol{v}_{2m})^{t_{1}}(I,\boldsymbol{v}_{2m+1})^{t_{2}}(I,\boldsymbol{w}_{2m})^{t_{3}}(I,\boldsymbol{w}_{2m+1})^{t_{4}}=(I,\boldsymbol{0}).$
@@
In all cases, $(I,\boldsymbol{0})$ can be represented by a full-image word, so $\langle\mathcal{G}\rangle$ is a group by Lemma 1.
@@ 5.6. $H$ is generated by a positive scale $A$
These are the forms of $A_{i}$ and $\boldsymbol{a}_{i}$ under the new basis $(\boldsymbol{x}_{V},\boldsymbol{x}_{W})$, where $\boldsymbol{x}_{V},\boldsymbol{x}_{W}$ are eigenvectors of $A$ such that $A\boldsymbol{x}_{V}=\lambda\boldsymbol{x}_{V},A\boldsymbol{x}_{W}=\lambda^{-1}\boldsymbol{x}_{W}$.
@@
$d_{ija}\coloneqq-\frac{a_{i}}{1-\lambda^{n_{i}}}+\frac{a_{j}}{1-\lambda^{-n_{j}}},\;\;d_{ijb}\coloneqq\frac{b_{i}}{1-\lambda^{-n_{i}}}-\frac{b_{j}}{1-\lambda^{n_{j}}}.$
@@
$({A^{\prime}_{i}},\boldsymbol{a}^{\prime}_{i})^{n_{j}}({A^{\prime}_{j}},\boldsymbol{a}^{\prime}_{j})^{n_{i}}=\left(I,(1-\lambda^{-n_{i}n_{j}})\cdot\left(\lambda^{n_{i}n_{j}}d_{ija},d_{ijb}\right)^{\top}\right)$
for all $(i,j)\in J_{+}\times J_{-}$.
@@
$({A^{\prime}_{k}},\boldsymbol{a}^{\prime}_{k})^{n_{k}}=\left(I,\boldsymbol{e}_{k}\right)$
for all $k\in J_{0}$. Denote by $\mathcal{G}^{\prime}$ the new alphabet $\{(A^{\prime}_{1},\boldsymbol{a}^{\prime}_{1}),\ldots,(A^{\prime}_{K},\boldsymbol{a}^{\prime}_{K})\}$. The following lemma shows that the rays $\boldsymbol{v}\mathbb{R}_{>0}$ such that $(I,\boldsymbol{v})\in\langle\mathcal{G}^{\prime}\rangle$ can reach both “extremities” of the cells $\operatorname{Cell}(\boldsymbol{d}_{ij}),(i,j)\in J_{+}\times J_{-}$. See Figure 10 for an illustration.
@@
Let $(i,j)$ be a pair in $J_{+}\times J_{-}$ and $(x,y)^{\top}\in\mathbb{R}^{2}$ be a vector in $\operatorname{Cell}(\boldsymbol{d}_{ij})$. If $\lambda>1$, then there exist elements $(I,\boldsymbol{v})$ and $(I,\boldsymbol{w})$ in $\langle\mathcal{G}^{\prime}\rangle$, where the vector $(x,y)^{\top}$ can be written as $r_{1}\boldsymbol{v}+r_{2}\boldsymbol{w}$ for some $r_{1},r_{2}\in\mathbb{R}_{>0}$. Furthermore, $(I,\boldsymbol{v})$ and $(I,\boldsymbol{w})$ are represented by words over the alphabet $\mathcal{G}^{\prime}$, such that the letters $(A^{\prime}_{i},\boldsymbol{a}^{\prime}_{i})$ and $(A^{\prime}_{j},\boldsymbol{a}^{\prime}_{j})$ both occur.
@@
For any $p\in\mathbb{Z}_{>0}$, let $\boldsymbol{v}_{p},\boldsymbol{w}_{p}$ be such that
@@
If $\operatorname{Cell}(\boldsymbol{d}_{ij})$ is of dimension one or zero, then $d_{ija}=0$ or $d_{ijb}=0$. In both cases, $(x,y)^{\top},\boldsymbol{v}_{p},\boldsymbol{w}_{p}$ are linearly dependent and have the same direction, so $(x,y)^{\top}$ can be written as $r_{1}\boldsymbol{v}_{p}+r_{2}\boldsymbol{w}_{p}$ for some $r_{1},r_{2}\in\mathbb{R}_{>0}$.
If $\operatorname{Cell}(\boldsymbol{d}_{ij})$ is of dimension two, then $d_{ija}\neq 0$ and $d_{ijb}\neq 0$. Let $p$ be large enough so that $\frac{x}{y}$ is between $\frac{\lambda^{pn_{i}n_{j}}d_{ija}}{d_{ijb}}$ and $\frac{d_{ija}}{\lambda^{pn_{i}n_{j}}d_{ijb}}$. Then $(x,y)^{\top}$ can be written as $r_{1}\boldsymbol{v}_{p}+r_{2}\boldsymbol{w}_{p}$ for some $r_{1},r_{2}\in\mathbb{R}_{>0}$.
@@
Suppose $J_{+}$ and $J_{-}$ are non-empty. Let $k\in J_{0}$, and $(x,y)^{\top}\in\operatorname{Cell}(\boldsymbol{e}_{k})$. If $\lambda>1$, then there exist elements $(I,\boldsymbol{v})$ and $(I,\boldsymbol{w})$ in $\langle\mathcal{G}^{\prime}\rangle$, such that the vector $(x,y)^{\top}$ can be written as $r_{1}\boldsymbol{v}+r_{2}\boldsymbol{w}$ for some $r_{1},r_{2}\in\mathbb{R}_{>0}$. Furthermore, $(I,\boldsymbol{v})$ and $(I,\boldsymbol{w})$ are represented as words over the alphabet $\mathcal{G}^{\prime}$, such that the letter $(A^{\prime}_{k},\boldsymbol{a}^{\prime}_{k})$ occurs.
@@
If $\operatorname{Cell}(\boldsymbol{e}_{k})$ has dimension one or zero, then take $p=0$ and the statement is trivial. Suppose that $\operatorname{Cell}(\boldsymbol{e}_{k})$ has dimension two, then we have $a_{k}\neq 0$ and $b_{k}\neq 0$. Fix a large enough $p$ so that $\frac{x}{y}$ is between $\frac{\lambda^{pn_{i}n_{j}}a_{k}}{\lambda^{-pn_{i}n_{j}}b_{k}}$ and $\frac{\lambda^{-pn_{i}n_{j}}a_{k}}{\lambda^{pn_{i}n_{j}}b_{k}}$. Then $(x,y)^{\top}$ can be written as $r_{1}(\lambda^{pn_{i}n_{j}}a_{k},\lambda^{-pn_{i}n_{j}}b_{k})^{\top}+r_{2}(\lambda^{-pn_{i}n_{j}}a_{k},\lambda^{pn_{i}n_{j}}b_{k})^{\top}$ for some $r_{1},r_{2}\in\mathbb{R}_{>0}$. When $q$ tends towards infinity, the direction of the vectors $\boldsymbol{v}_{pq},\boldsymbol{w}_{pq}$ tends respectively to $(\lambda^{pn_{i}n_{j}}a_{k},\lambda^{-pn_{i}n_{j}}b_{k})^{\top}$ and $(\lambda^{-pn_{i}n_{j}}a_{k},\lambda^{pn_{i}n_{j}}b_{k})^{\top}$. Therefore, for large enough $q$, the vector $(x,y)^{\top}$ can be written as $r^{\prime}_{1}\boldsymbol{v}_{pq}+r^{\prime}_{2}\boldsymbol{w}_{pq}$ for some $r^{\prime}_{1},r^{\prime}_{2}\in\mathbb{R}_{>0}$.
@@
  1. (1)
    $t_{1}t_{M}=1$. Suppose $t_{1}=t_{M}=1$, the case where $t_{1}=t_{M}=-1$ can be done analogously.
@@
    where $\boldsymbol{y}$ and $\boldsymbol{y}^{\prime}$ can be written as an $\mathbb{R}_{>0}$-linear combination of elements in $\mathcal{S}$. Therefore $\boldsymbol{x}=\boldsymbol{y}+\boldsymbol{y}^{\prime}$ also satisfies this claim.
  2. (2)
    $t_{1}t_{M}=-1$. In this case, $C_{2}\cdots C_{M-1}=C_{1}C_{2}\cdots C_{M}$ $=I$. Define
@@
    Then the product of $w^{\prime}$ is of the form $(I,\boldsymbol{x}^{\prime})$, where $\boldsymbol{x}^{\prime}$ is either $\boldsymbol{0}$ (when $w^{\prime}$ is empty), or $\boldsymbol{x}^{\prime}$ is an $\mathbb{R}_{>0}$-combination of elements in $\mathcal{S}$ by the induction hypothesis. Hence